I need help in solving this equation
solve if possible...
8 - 9t = 21t - 17
I get
- 9t = -9
-t = 1

Answer by edjones(8007)

8-9t=21t-17
25=30t add 9t and 17 to both sides.
t=25/30=5/6 divide each side by 30.
Check:
8-9*(5/6)=21*(5/6)-17
8-45/6=105/6-17
48-45=105-102 multiply both sides times 6.
3=3
Ed